OpenStudy (anonymous):

give exact and approximate solutions to three decimal place (x-8)^2=9

OpenStudy (anonymous):

x-8 = +3 or x-8 = -3 the answers are integers

OpenStudy (anonymous):

\[(x-8)^2=9\]\[x-8=\pm3\]\[x=\pm3+8\]\[x=5\]or\[x=11\]